OpenStudy (anonymous):

simplify x+7/x^2+4x-21

OpenStudy (saifoo.khan):

Factor the denominator.

OpenStudy (anonymous):

O great saifoo, how do you factor ?

OpenStudy (saifoo.khan):

Do you know how to factor? @swimmergirl

OpenStudy (mertsj):

Look at the numerator. That gives you a clue as to what one factor of the denominator might be and then you can figure out the other factor.

OpenStudy (anonymous):

-4x^3+21x+147/21x^2

OpenStudy (anonymous):

oh gosh I don't know how to factor....

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=X7B_tH4O-_s

OpenStudy (callisto):

Factorise x^2+4x-21: x^2+4x-21 = x^2 -3x + 7x -21 = x(x-3) + 7(x-3) = (x+7)(x-3) Do you understand?

OpenStudy (anonymous):